A satellite is spinning at a rate of 0.2 rev/s. The satellite has a main body that is spherical, with a mass of 586 kg and radius 2.4 m. What is the angular momentum of the satellite? Use four sig figs.
Answer #1

Then angular momentum is given by, L=Iw

Where I is the moment of inertia of the spherical satellite about the rotation axis and w is the angular speed.

Now, I=(2/5)mr2, where m is the mass of the satellite (=586kg) and r is the radius (=2.4m).

Thus, I ==mr? =* * 586 * (2.4)2 = 1350.14

And, w = 0.2rev/s = (0.2 * 27 )rad/s = 1.26rad/s

Therefore, angular momentum,

L = lw = 1350.14 * 1.26 = 1701.1764kg.m/s